-4/Y-1+4/Y+1=-8/y^2-1
Answer by JoelSchwartz(130)

y^2(-4/y+4/y=-8/y^2-1)
-4y+4y=-8-y^2
0=-8-y^2
8=-y^2
2*(2)^(1/2)=-y
y=-(8)^(1/2)